Solve the following systems by using either addition or substitution. If a unique solution does not exist, state whether the system is dependent or inconsistent.
10x+2y=7
y=-5x+3

Answer by rapaljer(4671)

Substitute y from the second equation into the first equation:
10x+2y=7
y=-5x+3

10x +2(-5x+3) = 7
10x -10x + 6=7
6=7

This statement is NEVER true, so there is NO SOLUTION. This means that the lines represented by these equations are parallel. The system is said to be "inconsistent".

R^2